# The Illusion of Diminishing Returns: Measuring Long Horizon Execution in LLMs

### v2

@@ 2.1 Diminishing Returns in Step Accuracy Compound Over a Long Horizon
We begin by analyzing the relationship between a model’s single-step accuracy and its horizon length. Note that this analysis applies not just to execution, but rather to any general long-horizon task. To obtain a mathematical relation, we make two simplifying assumptions similar to LeCun (2023). First, we assume a model’s step accuracy remains constant over the task. Second, we assume a model does not self-correct, meaning any single error leads to task failure. We assume this only for the analysis here, which is illustrative and provides useful intuition. Our subsequent empirical analysis goes beyond this, investigating how LLMs, in fact, do not exhibit constant step accuracy for long-horizon execution, and may also correct their mistakes.
@@
Assuming a constant step accuracy $p$ and no self-correction, the task-length $H$ at which a model achieves a success rate $s$ is given by:
@@
(The derivation is provided in Appendix˜H.)
@@ 3 Experiments
We choose short English words and two-digit integers to minimize errors arising from tokenization. We provide few-shot examples to clarify the task. More details, including the exact prompt, are provided in Appendix˜E. We also disentangle performance on the individual retrieval and composition operations, finding that models have much higher accuracies on each of them alone (Appendix˜D).
@@ 3.1 Effect of increasing the number of turns
Setup. We evaluate the Qwen3 (Yang et al., 2025) and Gemma3 (Gemma-Team et al., 2025) model families, as they offer a range of sizes: [4, 8, 14, 32]B and [4, 12, 27]B parameters, respectively. For this experiment, we set the turn complexity to its simplest form ($K=1$), providing a single key per turn, and vary the number of turns. Models are instructed to output the final answer directly, without intermediate thinking tokens, with the format enforced via few-shot examples. We verify that format-following errors are not the primary failure mode (Appendix˜F). We also show that the results below hold with chain-of-thought (CoT) prompting, and thinking models (Appendix Figure˜12), and the trends are not affected by the temperature used (Appendix Figure˜13).
@@
Result 2: Non-diminishing benefits of scaling model size. As shown in Figure˜4 (c), larger models sustain higher task accuracy for significantly more turns, resulting in a clear scaling trend for horizon length (Figure˜4 (b)). We abstain from deriving a “scaling law” since we can only obtain at most four model sizes from the same family, but the improvements do not seem diminishing. This observation is non-trivial. While the benefits of increasing model size are often attributed to improved capacity for knowledge, our task is not knowledge-constrained, as models achieve near-perfect first step accuracy (Figure˜4 (a)), nor is the task more complex so that a larger model would be required. Yet, larger models are clearly more reliable at executing the task for longer. A possible explanation is the redundancy of internal circuits in larger models, which ensembles to reduce error (Lindsey et al., 2025). However, we find that simulating this redundancy with output-level aggregation of parallel compute does not replicate the gains observed from scaling model size (Appendix˜B).
@@ 3.2 Why Does Turn Accuracy Degrade? The Self-Conditioning Effect
1. Degradation as the context length increases. The model’s performance degrades simply due to increasing context length (Zhou et al., 2025a), irrespective of its content.
@@
Result 3: Self-conditioning causes degradation in turn accuracy beyond long-context. Our results in Figure˜5 (a) show evidence for degradation due to both long-context and self-conditioning. When conditioned on an error-free history (Induced Error Rate = 0.00), model turn accuracy at turn 100 is below its initial value, consistent with prior observations of long-context degradation (Zhou et al., 2025a). More interestingly, as we increase the rate of injected errors into the context, accuracy at turn 100 consistently degrades further. This demonstrates the self-conditioning effect—as models make mistakes, they become more likely to make more mistakes, leading to a continuous degradation in per-turn accuracy throughout the output trajectory as shown in Figure˜5 (b).
Result 4: Unlike long-context, scaling model size does not mitigate self-conditioning. At the error rate of 0%, notice that the accuracy at turn 100 consistently improves as you scale model size. As shown in Figure˜5 (b), scaling to frontier (200B+ parameter) models like Kimi-K2 (Kimi-Team et al., 2025), DeepSeek-V3 (DeepSeek-AI et al., 2025), and Qwen3-235B-Instruct-2507 (Yang et al., 2025) largely solves long-context degradation for up to 100 turns, achieving near-perfect accuracy on a healed history. However, even these large models remain susceptible to self-conditioning, as their performance consistently degrades as the induced error rate in their history increases. This may be akin to recent results showing larger models shift in personality during multi-turn conversations (Choi et al., 2024; Becker et al., 2025), where in our case, the drift is toward a personality that makes errors.
@@
In Appendix˜G, we try the above setup of output manipulations with CoT prompting, finding that accuracy still deteriorates as the induced error rate increases. A potential confounder is that the manipulated outputs deviate from the CoT. We try to mitigate this issue with programmatically generated CoT traces, but still observe self-conditioning. We also try removing CoT traces from previous turns from history, which causes the model to stop using CoT completely. So, we now present results for the Qwen3 thinking models, which are trained with reinforcement learning (RL) to think even when previous turn traces are not presented. As before, we observe their turn 100 accuracy while controlling the error rate in prior turns.
@@
Result 5: Thinking fixes self-conditioning. In Figure˜6, we observe that the Qwen3 thinking models do not self-condition—the accuracy of the models at turn 100 remains stable, regardless of the error rate in its context. This could arise from two reasons. First, RL training can reduce the most likely next token prediction behaviour of language models, making them oriented towards task success rather than continuing the context. Second, the removal of thinking traces from prior turns could reduce the influence of prior turns on the model’s output, as it thinks about the new turn independently. By inspecting the models’ thinking traces, we observe that they do not refer back to their answers in prior turns, which could be a potential reason why they do not self-condition. Inspired by this, for non-thinking models, we experiment with context engineering by explicitly removing prior history and find that it indeed mitigates self-conditioning (Section˜A.2). We also find that just prompting models to self-verify their answers does not solve self-conditioning completely (Section˜A.1).
@@ 3.3 What is the length of tasks models can complete in a single turn?
In the previous sections, we measured how many turns models can successfully execute a single retrieve-then-compose step. However, most real-world tasks require more complex processing every turn. In fact, in Appendix˜C we show that the horizon length of different models can vary significantly at different turn complexities. The total task length a model can handle is a function of both the number of turns and the number of steps to execute per turn. We now measure the latter dimension: the maximum number of steps a model can execute per turn.
@@
Result 6: Without CoT, non-thinking models struggle to chain more than a few steps per turn. In Figure˜7 (left), we first find that when prompted to answer directly, without chain-of-thought, the larger Qwen3 32B, Gemma3 27B, as well as frontier non-thinking models like DeepSeek-V3 (670B), and Kimi K2 (1026B), fail to execute even a turn complexity of more than six. This is consistent with prior work showing the necessity of thinking tokens for transformers to perform sequential tasks (Weiss et al., 2021; Merrill and Sabharwal, 2023). We see that the number of steps the model can execute in a single turn improves significantly with chain-of-thought. This reinforces the importance of reasoning before acting (ReAct (Yao et al., 2023)) for agents, even if this costs more and fills up the context window. We show preliminary evidence that parallel test time compute is not as helpful, with majority voting leading to only marginal improvements in execution length (Appendix˜B).
Result 7: Benchmarking frontier models. In Figure˜7 (right), we benchmark frontier models on the length of task they can execute in a single turn. We find a surprisingly large gap between GPT-5 (codenamed Horizon) with 2176 steps and others like Claude-4 Sonnet (432 steps), Grok 4 (384 steps), and Gemini 2.5 Pro (120 steps). Overall, even our simple task can separate frontier models in their long-horizon execution capability, and presents a clear opportunity to improve current open-weight models.
@@ 4 Related Work
Increasing Task Length. Multiple works have recently shown how models worsen as problem complexity increases (Zhou et al., 2025b), often attributed to failures of reasoning (Cheng, 2025; Shojaee et al., 2025). Recently, multiple real-world long-horizon agentic benchmarks have been proposed (Backlund and Petersson, 2025; Xie et al., 2024; Shen et al., 2025), where prior work has studied planning failures (Chen et al., 2024b). By designing a task where no reasoning is required, given that we provide the model the requisite plan and knowledge, we show that execution alone can be a challenge, degrading model accuracy on longer tasks. Our observations on scaling could hold for the related problem of length-generalization—training models to succeed on tasks longer than those seen during training (Fan et al., 2024; Cai et al., 2025).
Long Context. Much of prior work has focused on improving the maximum context length that can be provided in the input to a language model (Su et al., 2021), and evaluating whether (Tay et al., 2020) and how (Olsson et al., 2022; Li et al., 2023) models maintain performance as the context gets longer (Tay et al., 2020). Closest is the recent RULER (Hsieh et al., 2024) and GSM-Infinite (Zhou et al., 2025b), which also use synthetic data to systematically evaluate long-context abilities. While long-context will help models execute for longer, it is a different capability compared to long-horizon execution (Zhou et al., 2023; Chen et al., 2024a), as it focuses on performance as a function of input, not output length. We identified one such difference, the self-conditioning effect—where past errors in model output increase the chance of future mistakes, and disentangle this effect from long-context degradation in Section˜3.2. Motivated by a similar effect at the token level, termed exposure bias, Ranzato et al. (2015) proposed training language models with RL.
@@
Tool Use. In symbolic AI, once tasks are formalized, for example, into STRIPS plans (Fikes and Nilsson, 1971), they can be executed without issues. Prior work (Chen et al., 2024b; Valmeekam et al., 2024) has shown LLMs struggle to match symbolic algorithms for automated planning. In contrast, we show LLMs can fail on straightforward execution (Zhu et al., 2025; Sun et al., 2025; Stechly et al., 2024) over a long horizon even when the plan is provided. Teaching LLMs to use tools offers one way to shift the burden of execution from probabilistic models to reliable programs (Schick et al., 2023). However, reasoning is often fuzzy and not always easy to implement as a tool, requiring the model to execute some steps by itself. Even calling the right tools requires reliable execution from the model (Patil et al., 2025).
Reliability. While there has been recent interest in evaluations for LLM reliability (Vendrow et al., 2025; Yao et al., 2024), it does not focus on long-horizon outputs, where the context differs every step. For example,  Yao et al. (2024) focus on the $pass^{k}$ metric, which checks if the model makes a mistake when we sample $k$ generations. However, this keeps the input fixed, and at 0 temperature (deterministic sampling), becomes equivalent to $pass@1$. In long-horizon tasks, error compounds irrespective of temperature, as shown in Appendix Figure˜13.
@@ 5 Conclusion
Limitations. As with any “synthetic” task (Allen-Zhu, 2024; Poli et al., 2024; Chollet et al., 2024) used for a controlled study of LLM capabilities, there are a few limitations of our setup. It does not reflect complexities and sources of error arising in real agentic tasks with a large number of possible actions. In such settings, the number of actions and the accuracy of each action can both vary based on the plan, requiring more careful consideration. It would be interesting future work to study the self-conditioning effect when doing diverse actions instead of repeating the same ones. Our results are observations about pretrained LLMs, and not inherent properties of transformers, so they might change with task-specific finetuning. Improvement on our task is necessary, but not sufficient for long-horizon execution on real-world tasks. Finally, our current task accuracy metric does not account for self-correction. In tasks where mistakes are acceptable and easy to undo, self-correction is a promising direction to improve long-horizon execution.

### v3

@@ 2.1 Diminishing Returns in Step Accuracy Compound Over a Long Horizon
We begin by analyzing the relationship between a model’s single-step accuracy and its horizon length. Note that this analysis applies not just to execution, but rather to any general long-horizon task. To obtain a mathematical relation, we make two simplifying assumptions similar to LeCun (2023). First, we assume a model’s step accuracy remains independent and constant over the task. Second, we assume a model does not self-correct, meaning any single error leads to task failure. We assume this only for the analysis here, which is illustrative and provides useful intuition. Our subsequent empirical analysis goes beyond this, investigating how LLMs, in fact, do not exhibit constant step accuracy for long-horizon execution, and may also correct their mistakes. Intuitively, if the model succeeds at one step with probability $p$, it’s success probability after $t$ steps is $p^{t}$. This is often negatively viewed as errors compounding over a long task. But conversely, it also implies that a small improvement in step accuracy can lead to a large increase in horizon length after a certain (high) threshold of single-step accuracy is achieved.
@@
Assuming an independent and constant step accuracy $p$ and no self-correction, the task-length $H$ at which a model achieves a success rate $s$ is given by:
@@
(The derivation is provided in Appendix˜J.)
@@ 3 Experiments
We choose short English words and two-digit integers to minimize errors arising from tokenization. We provide few-shot examples to clarify the task. More details, including the exact prompt, are provided in Appendix˜G. We also disentangle performance on the individual retrieval and composition operations, finding that models have much higher accuracies on each of them alone (Appendix˜F).
@@ 3.1 Effect of increasing the number of turns
Setup. We evaluate the Qwen3 (Yang et al., 2025a) and Gemma3 (Gemma-Team et al., 2025) model families, as they offer a range of sizes: [4, 8, 14, 32]B and [4, 12, 27]B parameters, respectively. For this experiment, we set the turn complexity to its simplest form ($K=1$), providing a single key per turn, and vary the number of turns. Models are instructed to output the final answer directly, without intermediate thinking tokens, with the format enforced via few-shot examples. We verify that format-following errors are not the primary failure mode (Appendix˜H). We also show that the results below hold with chain-of-thought (CoT) prompting, and thinking models (Appendix Figure˜13), and the trends are not affected by the temperature used (Appendix Figure˜14).
@@
Result 2: Non-diminishing benefits of scaling model size. As shown in Figure˜4 (c), larger models sustain higher task accuracy for significantly more turns, resulting in a clear scaling trend for horizon length (Figure˜4 (b)). We abstain from deriving a “scaling law” since we can only obtain at most four model sizes from the same family, but the improvements do not seem diminishing. This observation is non-trivial. While the benefits of increasing model size are often attributed to improved capacity for knowledge, our task is not knowledge-constrained, as models achieve near-perfect first step accuracy (Figure˜4 (a)), nor is the task more complex so that a larger model would be required. Yet, larger models are clearly more reliable at executing the task for longer. A possible explanation is the redundancy of internal circuits in larger models, which ensembles to reduce error (Lindsey et al., 2025). However, we find that simulating this redundancy with output-level aggregation of parallel compute does not replicate the gains observed from scaling model size (Appendix˜D).
@@ 3.2 Why Does Turn Accuracy Degrade? The Self-Conditioning Effect
1. Degradation as the context length increases. The model’s performance degrades simply due to increasing context length (Zhou et al., 2025b), irrespective of its content.
@@
Result 3: Self-conditioning causes degradation in turn accuracy beyond long-context. Our results in Figure˜5 (a) show evidence for degradation due to both long-context and self-conditioning. When conditioned on an error-free history (Induced Error Rate = 0.00), model turn accuracy at turn 100 is below its initial value, consistent with prior observations of long-context degradation (Zhou et al., 2025b). More interestingly, as we increase the rate of injected errors into the context, accuracy at turn 100 consistently degrades further. This demonstrates the self-conditioning effect—as models make mistakes, they become more likely to make more mistakes, leading to a continuous degradation in per-turn accuracy throughout the output trajectory as shown in Figure˜5 (b).
Result 4: Unlike long-context, scaling model size does not mitigate self-conditioning. At the error rate of 0%, notice that the accuracy at turn 100 consistently improves as you scale model size. As shown in Figure˜5 (b), scaling to frontier (200B+ parameter) models like Kimi-K2 (Kimi-Team et al., 2025), DeepSeek-V3 (DeepSeek-AI et al., 2025), and Qwen3-235B-Instruct-2507 (Yang et al., 2025a) largely solves long-context degradation for up to 100 turns, achieving near-perfect accuracy on a healed history. However, even these large models remain susceptible to self-conditioning, as their performance consistently degrades as the induced error rate in their history increases. A possible explanation for this is the relation between self-conditioning and in-context learning. In in-context learning, the model desirably conditions on user and environment inputs, for example, the correct few-shot demonstrations of the task we also provide in the initial prompt. In contrast, in self-conditioning, the model conditions on its own past behaviour, which could be wrong, leading to more mistakes. At a fundamental level, both arise from the model trying to output the most likely completion to its context, which larger models are known to be better at (Arora et al., 2025). Self-conditioning is also consistent with results showing how hallucinations snowball (Zhang et al., 2023), and larger models shift more in personality during multi-turn conversations (Choi et al., 2024; Becker et al., 2025), where in our case, the drift is toward a personality that makes errors.
@@
In Appendix˜I, we try the above setup of output manipulations with CoT prompting, finding that accuracy still deteriorates as the induced error rate increases. A potential confounder is that the manipulated outputs deviate from the CoT. We try to mitigate this issue with programmatically generated CoT traces, but still observe self-conditioning. We also try removing CoT traces from previous turns from history, which causes the model to stop using CoT completely. So, we now present results for the Qwen3 thinking models, which are trained with reinforcement learning (RL) to think even when previous turn traces are not presented. As before, we observe their turn 100 accuracy while controlling the error rate in prior turns.
Result 5: Thinking fixes self-conditioning. In Figure˜6, we observe that the Qwen3 thinking models do not self-condition—the accuracy of the models at turn 100 remains stable, regardless of the error rate in its context. This could arise from two reasons. First, RL training can reduce the most likely next token prediction behaviour of language models, making them oriented towards task success rather than continuing the context. Second, the removal of thinking traces from prior turns could reduce the influence of prior turns on the model’s output, as it thinks about the new turn independently. By inspecting the models’ thinking traces, we observe that they do not refer back to their answers in prior turns, which could be a potential reason why they do not self-condition. Inspired by this, for non-thinking models, we experiment with context engineering by explicitly removing prior history and find that it indeed mitigates self-conditioning (Section˜C.2). We also find that just prompting models to self-verify their answers does not solve self-conditioning completely (Section˜C.1).
@@ 3.3 What is the length of tasks models can complete in a single turn?
In the previous sections, we measured how many turns models can successfully execute a single retrieve-then-compose step. However, most real-world tasks require more complex processing every turn. In fact, in Appendix˜E we show that the horizon length of different models can vary significantly at different turn complexities. The total task length a model can handle is a function of both the number of turns and the number of steps to execute per turn. We now measure the latter dimension: the maximum number of steps a model can execute per turn.
@@
Result 6: Without CoT, non-thinking models struggle to chain more than a few steps per turn. In Figure˜7 (left), we first find that when prompted to answer directly, without chain-of-thought, the larger Qwen3 32B, Gemma3 27B, as well as frontier non-thinking models like DeepSeek-V3 (670B), and Kimi K2 (1026B), fail to execute even a turn complexity of more than six. This is consistent with prior work showing the necessity of thinking tokens for transformers to perform sequential tasks (Weiss et al., 2021; Merrill and Sabharwal, 2023). We see that the number of steps the model can execute in a single turn improves significantly with chain-of-thought. This reinforces the importance of reasoning before acting (ReAct (Yao et al., 2023)) for agents, even if this costs more and fills up the context window. We show preliminary evidence that parallel test time compute is not as helpful, with majority voting leading to only marginal improvements in execution length (Appendix˜D).
Result 7: Benchmarking frontier models. In Figure˜7 (right), we benchmark frontier models on the length of task they can execute in a single turn. We find a surprisingly large gap between GPT-5 (codenamed Horizon) with 2176 steps and others like Claude-4 Sonnet (432 steps), Grok 4 (384 steps), and Gemini 2.5 Pro (120 steps). Qwen3-Next, which combines Gated DeltaNet (Yang et al., 2025b) with standard attention, designed explicitly for long-context tasks, outperforms even larger models that use only standard attention, hinting at the impact of model architecture for long-horizon tasks. Overall, even our simple task can separate frontier models in their long-horizon execution capability, and presents a clear opportunity to improve current open-weight models.
@@ 4 Related Work
Increasing Task Length. Multiple works have recently shown how models worsen as problem complexity increases (Zhou et al., 2025a), often attributed to failures of reasoning (Cheng, 2025; Shojaee et al., 2025). Recently, multiple real-world long-horizon agentic benchmarks have been proposed (Backlund and Petersson, 2025; Xie et al., 2024; Shen et al., 2025), where prior work has studied planning failures (Chen et al., 2024b). By designing a task where no reasoning is required, given that we provide the model the requisite plan and knowledge, we show that execution alone can be a challenge, degrading model accuracy on longer tasks. Our observations on scaling could hold for the related problem of length-generalization—training models to succeed on tasks longer than those seen during training (Fan et al., 2024; Cai et al., 2025).
Long Context. Much of prior work has focused on improving the maximum context length that can be provided in the input to a language model (Su et al., 2021), and evaluating whether (Tay et al., 2020) and how (Olsson et al., 2022; Li et al., 2023) models maintain performance as the context gets longer (Tay et al., 2020). Closest is the recent RULER (Hsieh et al., 2024) and GSM-Infinite (Zhou et al., 2025a), which also use synthetic data to systematically evaluate long-context abilities. While long-context will help models execute for longer, it is a different capability compared to long-horizon execution (Zhou et al., 2023; Chen et al., 2024a), as it focuses on performance as a function of input, not output length. We identified one such difference, the self-conditioning effect—where past errors in model output increase the chance of future mistakes, and disentangle this effect from long-context degradation in Section˜3.2.
Self-correction. To address an issue similar to self-conditioning at the token level called “exposure bias”, Ranzato et al. (2015) proposed using reinforcement learning to train language models. This trains the model to react to mistakes in its prior outputs by correcting them, instead of making more of them, an intuition that also guides more recent work on RL for self-correction (Kumar et al., 2024; Ma et al., 2025). As a training-free fix for our setting, we also try model-based self-verification (see Kamoi et al. (2024) for a survey) at each turn in Appendix C.1, which can be considered a form of self-refinement (Madaan et al., 2023) without having access to environment feedback (Gou et al., 2023).
@@
Tool Use. In symbolic AI, once tasks are formalized, for example, into STRIPS plans (Fikes and Nilsson, 1971), they can be executed without issues. Prior work (Chen et al., 2024b; Valmeekam et al., 2024) has shown LLMs struggle to match symbolic algorithms for automated planning. In contrast, we show LLMs can fail on straightforward execution (Zhu et al., 2025b; Sun et al., 2025; Stechly et al., 2024) over a long horizon even when the plan is provided. Teaching LLMs to use tools offers one way to shift the burden of execution from probabilistic models to reliable programs (Schick et al., 2023). However, reasoning is often fuzzy and not always easy to implement as a tool, requiring the model to execute some steps by itself. Even calling the right tools requires reliable execution from the model (Patil et al., 2025).
@@ 5 Conclusion
Limitations. As with any “synthetic” task (Allen-Zhu, 2024; Poli et al., 2024; Chollet et al., 2024) used for a controlled study of LLM capabilities, there are a few limitations of our setup. Improvement on our task is necessary, but not sufficient for long-horizon execution on real-world tasks. It does not reflect the complexities and sources of error arising in real agentic tasks with a large number of possible actions. In such settings, the number of actions and the accuracy of each action can both vary based on the plan, requiring more careful consideration. Another issue in more complex tasks could be the possibility of multiple distinct correct plans, such that even if we provide a plan to the model, it could deviate from it while being correct. This is not possible in our controlled study, as using the wrong keys would likely lead to a wrong outcome. In Appendix˜A, we demonstrate how self-conditioning is a significant failure mode for LLM agents even in realistic benchmarks like ALFWorld (Shridhar et al., 2021), GAIA (Mialon et al., 2024), and WebShop (Yao et al., 2022). Our results are observations about current LLMs, and not inherent properties of transformers, so they might change with task-specific finetuning. Particularly, the single-turn task in Section˜3.3 is in theory parallelizable, which we discuss further and propose pathways to future-proofing via more inherently sequential tasks in Appendix˜B. Finally, our current task accuracy metric does not account for self-correction. In tasks where mistakes are acceptable and easy to undo, self-correction is a promising direction to improve long-horizon execution.
@@
Reproducibility statement
To ensure the reproducibility of our findings, we have provided comprehensive details of our methodology, experiments, and theoretical results. The mathematical derivation and assumptions for Proposition 1 are detailed in Appendix˜J. All specifics of our experimental methodology are consolidated in Appendix˜G, which covers our synthetic task design and data generation (Section˜G.1), the exact prompts used for both standard and thinking models (Sections˜G.2 and G.3), model specifications and hyperparameters like temperature and top-p (Section˜G.4), and the computational resources used for the evaluations (Section˜G.5). Furthermore, we have included the complete source code, data generation scripts, and the exact datapoints used as supplementary material.